It is assumed that the heights of American male adults are normally distributed with mean 69...

  1. It is assumed that the heights of American male adults are normally distributed with mean 69 inches and standard deviation 2.5 inches. What is the cut off value for the tallest 2.5% of all heights?
  1. 73.9
  2. 73.125
  3. 64.1
  4. Other. Please specify: __________.
Answer #1

solutuon

Given that,

mean = = 69

standard deviation = =2.5

Using standard normal table,

P(Z > z) = 2.5%

= 1 - P(Z < z) = 0.025

= P(Z < z ) = 1 - 0.025

= P(Z < z ) = 0.975

z = 1.96 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 1.96 *2.5+69

x= 73.9